OpenStudy (anonymous):

Use the commutative property to rewrite the expression: negative 3 y squared + 5 y

OpenStudy (zpupster):

the commutative property of addition is a+b= b+a just write it out

OpenStudy (skullpatrol):

$$\huge - 3y^2+5$$

OpenStudy (anonymous):

-3y^2 + 5 = 5 +-3y^2 ?

OpenStudy (zpupster):

don't forget the y in 5y

OpenStudy (anonymous):

i knowits a dumb question but whts the answer ?

OpenStudy (zpupster):

you did it right but your question said 5y not just 5

OpenStudy (anonymous):

A. negative 5 y plus 3 y squared B. 2 y cubed C. negative 15 y cubed D. 5 y minus 3 y squared

OpenStudy (anonymous):

? @zpupster

OpenStudy (anonymous):

?